Crew Photos – January 4th

Good evening/morning CAMPCOM,

Apologies for the belated submission of our Daily Photos for SOL 1.

We were wrapping up our training, filing reports, and getting through outstanding tasks of our own, due prior to the start of our Sim in a few hours.

Here are our 7 photos from SOL 1, all taken in and around the Hab.

The Photo of the Day is captioned "SOL 1 was a Pre-Sim day at MDRS238".

We’re starting low-key, but since it’s day 1, we thought we’d send you a couple of freebies too — our official crew portraits in both "Explorer Mode" and "Cool Mode."

You may also have noticed our little stowaway, nabbed last night after a dinner-time raid of our staterooms. She was a little scared, but by the AM had grown fond of us (maybe?) and we had grown fond of her. We named her Alice, fed her some oatmeal for the road, then dropped her off beyond the property limits just before sunset. It was kinda cute.

Please let me know if the submission formats are not appropriate, and I’ll be happy to correct.

To be continued at sign-off tomorrow.


P.J. Marcellino


Operations Report – January 3rd

Crew 238 Operations Report 03-01-2022


Name of person filing report: Simon Werner

Non-nominal systems: 1 CO- Sensor, lower deck, Atila takes care for
replacing with new one

Notes on non-nominal systems: Sensor went to low battery alarm/ test
mode last night twice. Change of Batteries did not solve the problem.
Measured batteries voltages today with multimeter, all were fully
loaded, so the sensor itself is damaged, See separate mail to mission
control this morning.


Spirit rover used: Yes

Hours: (before EVA) not noted (introduction)

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 98%

Currently charging: yes

Opportunity rover used: yes

Hours: not noted (introduction)

Beginning charge: 100%

Ending charge: 99%

Currently charging: yes

Curiosity rover used: yes

Hours: not noted (introduction)

Beginning charge: 100%

Ending charge: 99%

Currently charging: yes

Perseverance rover used: no

Hours: unknown

Beginning charge: unknown

Ending charge: unknown

Currently charging: unknown

General notes and comments: none

Summary of Hab operations: We’ve had an introduction to the spacesuits
today, the crew members selected their assigned suits and headsets.
New toilet was delivered.


Water (static tank): 507

Water (loft tank): 37 gallons, filled to 55 gallons

Water Meter: pre pumping 01543756 units, post pumping 01543757

Static to Loft Pump used – yes

Static tank pipe heater (on or off): New heater will be delivered and
installed tomorrow.

Static tank heater (On or off) unknown

Toilet tank emptied: yes

Summary of internet: nominal

Summary of suits and radios: suits assigned to team members, used for
introduction, Spacesuits in spacesuits storage room are nominal,
(number 3 had a red-light power supply, meanwhile turned green), all
plugged in. Spacesuits firstfloor: all nominal

Summary of GreenHab operations:

WATER USE: 240 gallons (guess)

Heater: On

Supplemental light: On

Harvest: none

Summary of ScienceDome operations: none

Dual split: Heat or AC, On or Off: OFF

Summary of RAM operations: Took multimeter and screwdriver set to Hab,
checked batteries and CO-sensor, will be brought back tomorrow.

Summary of any observatory issues: astronomy box is missing, lock of
observatory might be frozen in (Atila and Aga tested), Aga gets in
touch with Peter Detterline

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support:

Please pick up trash. Mr. Clean, toilet paper, kitchen paper towels
and pasta (Spaghetti) is needed. Thank you

Copyright © The Mars Society. All rights reserved. | Main Site